OpenStudy (anonymous):

Find the General solution of\[tanx + \tan(x+60) + \tan(x+120) = 3\]

OpenStudy (raden):

i think we can use identity : tan(A+B) = (tanA +tanB)/(1-tanAtanB) to simplify of tan(x+60) and tan(x+120)

OpenStudy (anonymous):

But it will end up in tan^3 x - 3tan^2 x - 3tan x + 1=0

OpenStudy (raden):

yes, factor out to get (tanx+1)(tan^2 x -4tanx + 1) = 0

OpenStudy (anonymous):

got..it thaxxx

OpenStudy (raden):

for zeros, tanx = -1 so on ..... bla bla :)

OpenStudy (raden):

welcome